# On $L^{p}$ extremals for Fourier extension estimate to fractional surface

### v1

@@ Abstract.
This article investigates the Fourier extension operator associated to the fractional surface $(\xi,|\xi|^{\alpha})$ with $\alpha\geq 2$. We show that nearly all valid scale-invariant Fourier extension inequalities possess extremals. More precisely, if the Fourier extension operator is bounded for an endpoint $p_{0}$, then for all $1<p<p_{0}$, the corresponding $L^{p}$-extremal sequences are precompact up to symmetries. In particular, we have developed the relevant $L^{p}$-profile decomposition for these Fourier extension operators.
@@
This work was supported in part by the National Key R&D Program of China [Grant No. 2023YFC3007303], the National Natural Science Foundation of China [Grant Nos. 12071052 & 12271501] and the China Postdoctoral Science Foundation [Grant No. GZB20230812].
@@ 1. Introduction
For $\alpha\geq 2$, we introduce the fractional surface $\mathbb{P}^{d}_{\alpha}=\{(\xi,|\xi|^{\alpha}):\xi\in\mathbb{R}^{d}\}$ in $\mathbb{R}^{d+1}$, equipped with the following projection measure
@@
This serves as a natural generalization of the paraboloid, while one should further notice that this fractional surface can have zero Gaussian curvature at the origin. We investigate the associated Fourier extension operator given by
@@
In this setting, roughly speaking, the Fourier restriction theory investigates that whether $\mathcal{E}_{\alpha}$ is a bounded operator from $L^{p}(\mathbb{R}^{d})$ to $L^{q}(\mathbb{R}^{d+1})$ for some pair $(d,p,q)$. Indeed, applying some rescaling and stationary phase arguments [45, Chapter 8, page 363, Exercise 5.13], one conclude that the desired Fourier extension estimate
(1) $\|\mathcal{E}_{\alpha}f\|_{L^{q}(\mathbb{R}^{d+1})}\leq M_{\alpha,p}\|f\|_{L^{p}(\mathbb{R}^{d})}$
@@
The typical paraboloid case $\alpha=2$ has been studied by various mathematicians, and the well-known Fourier restriction conjecture for paraboloid predicts that the aforementioned necessary condition (2) is indeed sufficient. Up to now, this conjecture is solved for dimension $d=1$ and remains open for higher dimensions. For more background and progress on this Fourier restriction theory, the reader can see relevant papers such as [3, 4, 17, 24, 25, 26, 46, 48, 49, 50, 51, 52]. In addition, for the general fractional surface with $p=2$, the classical work of Kenig-Ponce-Vega [32] states that $\mathcal{E}_{\alpha}$ is a bounded operator from $L^{2}(\mathbb{R}^{d})$ to $L^{\frac{2d+2\alpha}{d}}(\mathbb{R}^{d+1})$. Note that this extension estimate corresponds to the Strichartz estimate for fractional Schrödinger equation with initial data in $L^{2}(\mathbb{R}^{d})$.
@@
Denote the group generated by these symmetries as $G$. Then a sequence of functions $\{f_{n}\}\subset L^{p}(\mathbb{R}^{d})$ is called precompact up to symmetries if there exists a sequence of symmetries $\{\mathcal{G}_{n}\}\subset G$ such that $\{\mathcal{G}_{n}f_{n}\}$ has a convergent subsequence in $L^{p}(\mathbb{R}^{d})$. For an extremal sequence, notice that the precompactness up to symmetries implies the existence of extremal function.
As mentioned above, there are abundant works investigating the extremal problems for Fourier extension operators. Here we briefly sketch some of the relevant works, and the reader can see the survey papers [21, 39] as well as the references therein for further details on recent progress.
  • Paraboloid with $p=2$: first for $d=1$, by applying some concentration-compactness arguments from Lions [36, 37], Kunze [33] showed the existence of extremals; then for $d=\{1,2\}$, by following different approaches, Foschi [19] and Hundertmark-Zharnitsky [28] independently showed that the only extremals are Gaussians modulo their symmetries; later for all dimension, by using the $L^{2}$-based profile decomposition from Bégout-Vargas [1], Shao [43] showed the existence of extremals.
  • Fractional surface with $p=2$: first for $(\alpha,d)=(4,1)$ and $(\alpha,d)=(4,d)$, by establishing the corresponding $L^{2}$-linear profile decompositions, Jiang-Pausader-Shao [29] and Jiang-Shao-Stovall [30] established the characterizations on precompactness of extremal sequences; then for $(\alpha,d)=(4,2)$ and $(\alpha,d)=(\alpha_{1},1)$ with $\alpha_{1}\in(1,\alpha_{0})$ for some $\alpha_{0}>4$, by applying some comparison principle for convolutions of certain singular measures, Oliveira e Silva-Quilodrán [40] and Brocchi-Oliveira e Silva-Quilodrán [5] resolved the aforementioned characterizations and obtained the existence of extremals; later for $(\alpha,d)=(\alpha,1)$ and arbitrary $(\alpha,d)$, by establishing the corresponding $L^{2}$-based profile decomposition and applying the missing mass method from Lieb [22, 34], the first and third authors [14, 13] established the characterization on precompactness of extremal sequences, which combining the result of [40, Proposition 6.9] can imply the existence of extremals when $(\alpha,d)=(\alpha_{2},2)$ with $\alpha_{2}\in(1,\alpha_{*})$ for some $\alpha_{*}>5$.
  • Other surfaces with $p=2$: such as the spheres [11, 20, 44, 22], cones [19, 6, 41], odd curves [5, 12, 23, 42], hyperboloids [7, 8], and non-endpoint type estimates [15, 16, 27].
It can be seen that all of these aforementioned results focus on the $L^{2}$-based Fourier extension inequalities. In fact, for paraboloid Fourier extension operator, the first $L^{p}$-based result is due to Christ-Quilodrán [10] in 2014, which showed that Gaussian functions are not extremals when $p\neq\{1,2\}$; then in 2020, Stovall [47] proved the precompactness of $L^{p}$-extremal sequences conditional on improvements toward the restriction conjecture; after that, by following the strategy given by Stovall, similar $L^{p}$-based results have been established for the moment curve [2], the sphere [18] and the cone [38].
In this article, we investigate the $L^{p}$-extremal theory for Fourier extension operators, where the associated surfaces have vanishing Gaussian curvature at some point. This phenomenon avoids us from directly applying Tao’s bilinear restriction estimate [48], which is a crucial tool for establishing the desired profile decomposition consequences. By following the strategy given by Stovall [47] and combining some previous conclusions of Di-Yan [13], for the fractional surface, we show that nearly all valid scale-invariant Fourier restriction inequalities have extremals. More precisely, our main result can be stated as follows.
@@
For $d\geq 2$ and $\alpha>2$, assume that $\mathcal{E}_{\alpha}$ extends as a bounded operator from $L^{p_{0}}(\mathbb{R}^{d})$ to $L^{q_{0}}(\mathbb{R}^{d+1})$ for some $1<p_{0}<q_{0}=\frac{d+\alpha}{d}p_{0}^{\prime}$. Then for the exponents $1<p<p_{0}$ with $q=\frac{d+\alpha}{d}p^{\prime}$, all extremal sequences for $M_{\alpha,p}$ are precompact up to symmetries, thus the extremal exists.
As mentioned above, the typical $\alpha=2$ case has been studied by Stovall [47] and the special $p=2$ case has been studied by many mathematicians. Here we pointed out that in some sense our assumption on $p_{0}$ is quite strong: noticing that $\mathcal{E}_{\alpha}$ is obviously $L^{1}(\mathbb{R}^{d})$ to $L^{\infty}(\mathbb{R}^{d+1})$ bounded, hence by interpolation, our assumption on $p_{0}$ implies that $\mathcal{E}_{\alpha}$ is bounded from $L^{p}(\mathbb{R}^{d})$ to $L^{q}(\mathbb{R}^{d+1})$ for all $1<p<p_{0}$ with $q=\frac{d+\alpha}{d}p^{\prime}$. This fact guarantees that we can investigate the $L^{p}$-extremal problems under the assumption on $p_{0}$.
To prove Theorem 1.1, we follow the outline given by Stovall [47]. Firstly, we establish some refined estimates to demonstrate that extremal sequences possess good frequency localization. Subsequently, the functions become bounded with compact support, which allows us to employ $L^{2}$-based profile decomposition to handle these functions. Here for the $L^{2}$-based profile decomposition, one technical problem arises from the fact that the fractional surface’s Gaussian curvature vanishing at the origin. Fortunately, the first and the second authors [13, Section 2] have addressed this problem: they divide the fractional surface into countable annuli, and then further divide these annuli by angles.
@@ 2.1. The $L^{2}$ consequences
First, we introduce a regular-type Fourier extension operator on fractional surface and some basic dyadic terminologies. We recall the following operator
@@
A dyadic interval is an interval of the form $[m2^{-n},(m+1)2^{-n}]$ with $m,n\in\mathbb{Z}$, and a dyadic cube is a product of dyadic intervals all having the same length; the set of all dyadic cubes of side length $2^{-k}$ is denoted by $\mathcal{D}_{k}$, and the set of all dyadic cubes is denoted by $\mathcal{D}$; for two dyadic cubes $\tau$ and $\tau^{\prime}$, the notation $\tau\sim\tau^{\prime}$ means that they have the same length $\ell(\tau)=\ell(\tau^{\prime})$ but their distance is like $C_{d,\alpha}\ell(\tau)$ with the constant $C_{d,\alpha}$ sufficiently large [^1] Further details can be seen in [13, Definition 2.4].
The well-known bilinear restriction estimate [48, Section 9, third remark] is also used in [13]. However, the Gaussian curvature of the fractional surface vanishes at the origin. To resolve this difficulty, one need to divide $\mathbb{R}^{d}$ into dyadic cube annuli as follows
@@
To utilize the Whitney decomposition, one need to further divide the dyadic cube annuli into several parts based on the angles [^2] Here the constant $K_{d,\alpha}$ is pretty large and further details can be seen in [13, Page 7]
@@
To illustrate the effect of frequency translation parameters and their asymptotic behavior, we recall some notations in [13, Section 4]. For $f\in L^{2}(\mathbb{R}^{d})$ and $\xi_{0}\in\mathbb{R}^{d}$, we introduce the operators
@@
Let $f\in L^{2}(\mathbb{R}^{d})$ with $\|f\|_{L^{2}(\mathbb{R}^{d})}=1$ and $\{\xi_{n}\}\subset\mathbb{R}^{d}$ with $|\xi_{n}|\to\infty$. Set the notations
@@
for sufficiently large $n$. As we will see later, this estimate (6) implies that a sequence of functions which looks like $\{f_{n}\}$ in Lemma 2.5 cannot be an extremal sequence for $M_{\alpha,p}$; on the other hand, based on the work of [43, Theorem 1.4] which claims the existence of extremals for $M_{2,2}$, this Lemma 2.5 immediately implies the fact $\widetilde{M}_{\alpha}\geq\mathbf{\mathrm{a}}^{*}_{d,\alpha}M_{2,2}$.
@@
Finally, by the endpoint Strichartz estimate of Keel-Tao [31, Theorem 1.2] and some stationary phase arguments, one can obtain the following Sobolev-type Strichartz estimate. Further details can be seen, for example, in the paper [9, Proof of Theorem 1.2].
@@ 2.2. The $L^{p}$ consequences
holds for arbitrary $L^{p}(X)$ functions $f$ and $g$. Then there exists $\theta_{0}\in(0,1)$ such that
@@
For the integers $j\in\mathbb{N}$ and $J\in\mathbb{N}$, define the following two operators
@@ 3. The $L^{2}$ theory
The $L^{2}$-based profile decomposition is one of the main ingredients to show the desired $L^{p}$-precompact result Theorem 1.1. Thus, in this section, we investigate the $L^{2}$ case first. As mentioned above, Di-Yan [13] have established the necessary preliminaries to establish the desired $L^{2}$-based profile decomposition. Here we will apply these preliminaries to show this $L^{2}$-decomposition in Subsection 3.1, and then give some corollaries in Subsection 3.2.
@@ 3.1. The $L^{2}$-based profile decomposition
By the classical Brézis-Lieb lemma [35, Section 1.9], we have
@@
Recall the inequality (6) which gives the desired dominating function in $L^{2+\frac{4}{d}}(\mathbb{R}^{d+1})$. Combining these three facts, a generalization of the Brézis-Lieb lemma [22, Lemma 3.1] implies
@@
(19) $\left(\frac{h_{n}^{1}}{h_{n}^{2}}\right)^{\frac{d}{2}}e^{i\left(h_{n}^{1}\left(x_{n}^{1}-x_{n}^{2}\right),\left(h_{n}^{1}\right)^{\alpha}\left(t_{n}^{1}-t_{n}^{2}\right)\right)\left(\cdot+\frac{\xi_{n}^{1}}{h_{n}^{1}},\left|\cdot+\frac{\xi_{n}^{1}}{h_{n}^{1}}\right|^{\alpha}\right)}\widetilde{g^{1}_{n}}\omega_{n}^{1}\left(\frac{h_{n}^{1}}{h_{n}^{2}}\cdot+\frac{\xi_{n}^{1}-\xi_{n}^{2}}{h_{n}^{2}}\right).$
@@ 3.2. Some applications of the profile decomposition
This subsection contains some corollaries of the profile decomposition obtained in the previous subsection.
First, combining the aforementioned profile decomposition and the asymptotic Schrödinger behavior Lemma 2.5, one can follow the classical arguments in [43, Proof of Theorem 1.4] to deduce the following characterization of precompactness for extremal sequences.
@@
Second, applying Sobolev embedding and stationary phase analysis, one can follow the arguments in [27, Proof of Theorem 2.4] to establish the following noncritical profile decomposition.
@@
Third, as this profile decomposition Corollary 3.5 does not contain the frequency translation parameters $\xi_{n}$, one can follow the arguments in [27, Proof of Theorem 1.2] to derive the following precompactness result.
@@
Let $\alpha>2$ and $d\geq 2$. Then all extremal sequences for $M_{\alpha,2}$ are precompact up to symmetries, thus the extremal exists.

### v2

@@ Abstract.
This article investigates the Fourier extension operator associated with the fractional surface $(\xi,|\xi|^{\alpha})$ for $\alpha\geq 2$. We show that the relevant $L^{p}\to L^{q}$ Fourier extension inequality possesses extremals for all exponents $p\in[1,2]$. Moreover, for all $p\in(1,2]$, the corresponding $L^{p}$-extremal sequences are precompact up to symmetries.
@@
This work was supported in part by the National Key R&D Program of China [Grant No. 2023YFC3007303], the National Natural Science Foundation of China [Grant Nos. 12071052 & 12271501] and the China Postdoctoral Science Foundation [Grant Nos. GZB20230812 & 2024M753436].
@@ 1. Introduction
For $\alpha\geq 2$, we investigate the fractional surface $\mathbb{P}^{d}_{\alpha}=\{(\xi,|\xi|^{\alpha}):\xi\in\mathbb{R}^{d}\}$ in $\mathbb{R}^{d+1}$, equipped with the following projection measure
@@
This serves as a natural generalization of the paraboloid, while one should notice that this fractional surface can have zero Gaussian curvature at the origin. The associated Fourier extension operator is given by
@@
In this setting, the Fourier restriction theory investigates whether $\mathcal{E}_{\alpha}$ is a bounded operator from $L^{p}(\mathbb{R}^{d})$ to $L^{q}(\mathbb{R}^{d+1})$ for some pair $(d,p,q)$. Indeed, applying some rescaling and stationary phase arguments [41, Chapter 8, page 363, Exercise 5.13], one concludes that the desired Fourier extension estimate
(1) $\|\mathcal{E}_{\alpha}f\|_{L^{q}(\mathbb{R}^{d+1})}\leq C_{\alpha,p}\|f\|_{L^{p}(\mathbb{R}^{d})}$
@@
The typical paraboloid case $\alpha=2$ is the well-known Fourier restriction conjecture for the paraboloid. Up to now, this conjecture has been solved for dimension $d=1$ and remains open for higher dimensions. For example, when $d=2$, this conjecture has been confirmed for $q>3+\frac{3}{14}$ by Wang-Wu [46]. For more related topics, one can refer to publications like [13, 45].
For general fractional surfaces with $\alpha\geq 2$, the relevant Fourier restriction phenomena have also been studied by several mathematicians, see for instance the papers [5, 27, 42]. Indeed, the classical work of Kenig-Ponce-Vega [27] states that $\mathcal{E}_{\alpha}$ is a bounded operator from $L^{2}(\mathbb{R}^{d})$ to $L^{\frac{2d+2\alpha}{d}}(\mathbb{R}^{d+1})$. Since the operator $\mathcal{E}_{\alpha}$ is obviously $L^{1}(\mathbb{R}^{d})$ to $L^{\infty}(\mathbb{R}^{d+1})$ bounded, by interpolation, one can see that $\mathcal{E}_{\alpha}$ is bounded from $L^{p}(\mathbb{R}^{d})$ to $L^{q}(\mathbb{R}^{d+1})$ for all $p\in[1,2]$ with $q=\frac{d+\alpha}{d}p^{\prime}$.
@@
Denote the group generated by these symmetries as $G$. Then a sequence of functions $\{f_{n}\}\subset L^{p}(\mathbb{R}^{d})$ is called precompact up to symmetries if there exists a sequence of symmetries $\{\mathcal{G}_{n}\}\subset G$ such that $\{\mathcal{G}_{n}f_{n}\}$ has a convergent subsequence in $L^{p}(\mathbb{R}^{d})$. For an extremal sequence, notice that the precompactness up to symmetries implies the existence of the extremal function.
For the case $p=2$, the extremal problems for Fourier extension operators have been well-studied. There are many approaches involved in these works such as concentration-compactness principle/missing mass method/profile decomposition argument [4, 11, 14, 20, 21, 25, 30, 38, 39, 40], Euler-Lagrange type equation [7, 12, 17, 18, 23, 24], geometric comparison principle [4, 36, 37], bilinear restriction theory [8, 15, 20, 25]. Indeed, the first and the third authors have summarized some of the relevant works in [14, 15], and the reader can also refer to the survey papers [19, 34, 35] as well as the references therein for further details on recent progress.
The first $L^{p}$-based result is due to Christ-Quilodrán [10] in 2014, who showed that Gaussian functions are not extremals when $p\neq\{1,2\}$, by investigating the associated Euler-Lagrange equation; then in 2020, Stovall [43] proved the precompactness of $L^{p}$-extremal sequences conditional on improvements toward the restriction conjecture, by generalizing the $L^{2}$-based profile decomposition argument in [39]; after that, the strategy given by Stovall is also used in investigating other surfaces or curves such as the moment curve [3], the sphere [16], the monomial curves [BS2024] and the cone [33].
In this article, we investigate the $L^{p}$-extremal theory for Fourier extension operators, where the associated surfaces (fractional surface) can have vanishing Gaussian curvature at some point. This vanishing-curvature property prevents us from directly applying Tao’s bilinear restriction estimate [44], which is a crucial tool for establishing the desired profile decomposition consequences. For the fractional surface, we obtain the following result.
@@
Let $d\geq 2$ and $\alpha>2$. Then for all $p\in[1,2]$, the extremals for $M_{\alpha,p}$ exist; moreover, for all $p\in(1,2]$, the extremal sequences for $M_{\alpha,p}$ are precompact up to symmetries.
Remark.
Recall that Kenig-Ponce-Vega’s result [27, Theorem 2.1] guarantees the boundedness of $M_{\alpha,p}$ for $p\in[1,2]$. In addition, if one can confirm that $M_{\alpha,p}$ is finite for $p\in[1,p_{0}]$ with some $p_{0}>2$, then our method will directly give the existence of extremals for $p\in[1,p_{0})$.
To prove Theorem 1.1, we follow the outline given by Stovall [43]. Firstly, one should establish some refined estimates to demonstrate that extremal sequences possess good frequency localization. Subsequently, the functions become bounded with compact support, which can be handled by employing the $L^{2}$-based profile decomposition. Here for establishing the $L^{2}$-based profile decomposition, one technical problem arises from the fact that the fractional surface’s Gaussian curvature vanishes at the origin. Fortunately, Di-Yan [15, Section 2] has addressed this problem: they divide the fractional surface into countable annuli, and then further divide each annuli into finite regions by dividing the angles. However, that paper [15] used a missing mass method rather than a profile decomposition argument. In this paper, to deal with the $L^{p}$ situation and follow Stovall’s outline, we will establish the full profile decomposition. Here we should remark that the profile decomposition consequence also has many applications in PDE, see for instance [1, 2, 6, 28, 29, 32].
@@ 2.1. The $L^{2}$ consequences
First, we introduce a regular-type Fourier extension operator on the fractional surface and some basic dyadic terminologies. We recall the following operator
@@
A dyadic interval is an interval of the form $[m2^{-n},(m+1)2^{-n}]$ with $m,n\in\mathbb{Z}$, and a dyadic cube is a product of dyadic intervals all having the same length; the set of all dyadic cubes of side length $2^{-k}$ is denoted by $\mathcal{D}_{k}$, and the set of all dyadic cubes is denoted by $\mathcal{D}$; for two dyadic cubes $\tau$ and $\tau^{\prime}$, the notation $\tau\sim\tau^{\prime}$ means that they have the same length $\ell(\tau)=\ell(\tau^{\prime})$ but their distance is like $C_{d,\alpha}\ell(\tau)$ with the constant $C_{d,\alpha}$ sufficiently large [^1] Further details can be seen in [15, Definition 2.4].
The well-known bilinear restriction estimate [44, Section 9, third remark] is also used in [15]. However, the Gaussian curvature of the fractional surface vanishes at the origin. To resolve this difficulty, one needs to divide $\mathbb{R}^{d}$ into dyadic cube annuli as follows
@@
To utilize the Whitney decomposition, one needs to further divide the dyadic cube annuli into several parts based on the angles [^2] Here the constant $K_{d,\alpha}$ is pretty large and further details can be seen in [15, Page 7]
@@
To illustrate the effect of frequency translation parameters and their asymptotic behavior, we recall some notations in [15, Section 4]. For $f\in L^{2}(\mathbb{R}^{d})$ and $\xi_{0}\in\mathbb{R}^{d}$, we introduce the operators
@@
Let $f\in L^{2}(\mathbb{R}^{d})$ with $\|f\|_{L^{2}(\mathbb{R}^{d})}=1$ and $\{\xi_{n}\}\subset\mathbb{R}^{d}$ with $|\xi_{n}|\to\infty$ as $n\to\infty$. Set the notations
@@
for sufficiently large $n$. As we will see later, this estimate (6) implies that a sequence of functions which looks like $\{f_{n}\}$ in Lemma 2.5 cannot be an extremal sequence for $M_{\alpha,p}$; on the other hand, based on the work of [39, Theorem 1.4] which claims the existence of extremals for $M_{2,2}$, this Lemma 2.5 immediately implies the fact $\widetilde{M}_{\alpha}\geq\mathbf{\mathrm{a}}^{*}_{d,\alpha}M_{2,2}$.
@@
Finally, by the endpoint Strichartz estimate of Keel-Tao [26, Theorem 1.2] and some stationary phase arguments, one can obtain the following Sobolev-type Strichartz estimate. Further details can be seen, for example, in the paper [9, Proof of Theorem 1.2].
@@ 2.2. The $L^{p}$ consequences
holds for arbitrary $L^{p}(X)$ functions $f$ and $g$. Then there exists $\theta\in(0,1)$ such that
@@
For the integers $j\in\mathbb{N}$ and $J\in\mathbb{N}$, define the following two projection-type operators
@@ 3. The $L^{2}$ theory
The $L^{2}$-based profile decomposition is one of the main ingredients to show the desired $L^{p}$-precompact result Theorem 1.1. Thus, in this section, we investigate the $L^{2}$ case first. As mentioned above, Di-Yan [15] has investigated the $L^{2}$-precompactness by the missing mass method and established some preliminaries. Here we will show the desired $L^{2}$-based profile decomposition in Subsection 3.1, and then give some corollaries in Subsection 3.2.
@@ 3.1. The $L^{2}$-based profile decomposition
By the classical Brézis-Lieb lemma [31, Section 1.9], we have
@@
Recall the inequality (6) which gives the desired dominating function in $L^{2+\frac{4}{d}}(\mathbb{R}^{d+1})$. Combining these three facts, a generalization of the Brézis-Lieb lemma [20, Lemma 3.1] implies
@@
(19) $\left(\frac{h_{n}^{1}}{h_{n}^{2}}\right)^{\frac{d}{2}}e^{i\left(h_{n}^{1}\left(x_{n}^{1}-x_{n}^{2}\right),\left(h_{n}^{1}\right)^{\alpha}\left(t_{n}^{1}-t_{n}^{2}\right)\right)\left(\cdot+\frac{\xi_{n}^{1}}{h_{n}^{1}},\left|\cdot+\frac{\xi_{n}^{1}}{h_{n}^{1}}\right|^{\alpha}\right)}\widetilde{g^{2}_{n}}\omega_{n}^{1}\left(\frac{h_{n}^{1}}{h_{n}^{2}}\cdot+\frac{\xi_{n}^{1}-\xi_{n}^{2}}{h_{n}^{2}}\right).$
@@ 3.2. Some applications of the profile decomposition
This subsection contains some corollaries of the profile decomposition Proposition 3.3 obtained in the previous subsection.
First, combining the aforementioned profile decomposition and the asymptotic Schrödinger behavior Lemma 2.5, one can follow the classical arguments in [39, Proof of Theorem 1.4] to deduce the following characterization of precompactness for extremal sequences.
@@
Second, applying the Sobolev embedding and stationary phase analysis, one can follow the arguments in [22, Proof of Theorem 2.4] to establish the following noncritical profile decomposition.
@@
Third, as this profile decomposition Corollary 3.5 does not contain the frequency translation parameters $\xi_{n}$, one can follow the arguments in [22, Proof of Theorem 1.2] to derive the following precompactness result.
@@
Let $\alpha>2$ and $d\geq 2$. Then all extremal sequences for $M_{\alpha,2}$ are precompact up to symmetries, thus extremals exist.